According to a local survey performed in January, 55% of the people polled said they would reelect the town's mayor. However, the mayor's percentage of votes decreased 2% each month. The survey revealed that 30% of the pople polled would vote for the challenger. The challengers percentage of votes increased 3% each month. In what month will the percentage of pople willing to vote for the challenger equal that of the mayor? What percentage will vote for the challenger in that month?
Answer by stanbon(75887)

According to a local survey performed in January, 55% of the people polled said they would reelect the town's mayor. However, the mayor's percentage of votes decreased 2% each month. The survey revealed that 30% of the pople polled would vote for the challenger. The challengers percentage of votes increased 3% each month. In what month will the percentage of pople willing to vote for the challenger equal that of the mayor? What percentage will vote for the challenger in that month?
----
Mayor's support:: M(x) = 55-2x
-----
Challenger support:: C(x) = 30+3x
----
Equation:
C(x) = M(x)
30+3x = 55-2x
5x = 25
x = 5 months
Ans: June
------------
Cheers,
Stan H.
-------